# Bridging the Sim2Real Gap: Vision Encoder Pre-Training for Visuomotor Policy Transfer

### v1

@@ IV-A Analyzing Domain Invariance - Action Inference Plot
This section outlines the results of evaluating the 23 surveyed encoders using our proposed quantitative metrics. In Figure 5, we plot the results of the centroid distance metric against the linear probing metric for each encoder to visualize their relative performance in domain invariance and action inference. CNN-based encoders are plotted in blue, while transformer-based encoders are in orange. The shape of each point indicates its form of pre-training: a circle for encoders trained on manipulation data and a square for encoders trained on general visual data. The relative size of each point represents the number of parameters in the corresponding encoder. The y-axis represents the action score, while the x-axis represents the domain invariance score, meaning being closer to 1 on both axes indicates higher performance in each metric, or that being further up and right on the plot is ideal for any given encoder.
@@
Figure 6: Grad-CAM and t-SNE Plot Generated With DinoV2-B
Figure 7: Grad-CAM and t-SNE Plot Generated With MCR
Figure 8: Grad-CAM and t-SNE Plot Generated With VGG-16
Figure 9: Grad-CAM and t-SNE Plot Generated With CLIP-Base-32
Figure 5 demonstrates general trends regarding a given encoder’s characteristics and performance across the two metrics. A significant pattern was the relationship between pre-training data and action-inference performance. All encoders trained on manipulation data performed highly in action inference, while encoders trained on general data had mixed results. This pattern is exhibited when comparing the ResNet models (pre-trained only on ImageNet [35]) to their R3M counterparts, where R3M encoders are ResNet-based but trained with self-supervised learning on vast amounts of publicly available manipulation data. Each R3M encoder outperformed its ResNet counterpart in both metrics. Another intriguing pattern was that, despite the general understanding that more model parameters produce more accurate results, the number of parameters didn’t seem to impact either performance metric, indicating that it doesn’t affect Sim2Real performance. As for encoder type, ViT-based encoders generally performed worse in the centroid distance metric than CNN-based encoders. According to the plot, the overall highest-performing encoder was MCR (numbered 8 on the plot) [22], a CNN-based encoder trained on manipulation data, abiding by the observed patterns. MCR is especially unique compared to other encoders because it is specifically pre-trained on robotic manipulation data, rather than human manipulation data.
IV-B Comparing Quantitative Results To Qualitative Observations
We analyze the results of our qualitative experimentation using Grad-CAM saliency maps and t-SNE plots to contextualize the performance of various encoders on the plot. For each encoder, we generated a grid of saliency maps and a t-SNE plot to observe how they relate to the encoder’s quantitative domain invariance and action inference performance. Figures 7 through 10 show the Grad-CAM maps and t-SNE plots generated from four selected models, including the overall worst-performing model (DinoV2-B), the overall best-performing model (MCR), a model with high domain invariance but low action inference ability (VGG-16), and a model with high action inference ability but low domain invariance (CLIP-Base-32).
Figure 6 shows that DinoV2-B does not have a clear focus across any of the maps, indicating low action inference performance, which corresponds with its action probe metric. Furthermore, the t-SNE plot for DinoV2-B shows a wide gap between the sim and real clusters, indicating low domain invariance, which its domain invariance score confirms.
Figure 7 shows that MCR not only has a sharp focus on the robot end-effectors and the objects in the Grad-CAM maps for both simulated and real images but also exhibits some cluster overlap in the t-SNE plot, being one of the only models to do so. This finding aligns well with its high performance in both quantitative metrics.
In Figure 8, VGG-16 demonstrates a lack of focus on any particular part of the images in the Grad-CAM maps but has a relatively small gap between sim and real clusters in the t-SNE plot, explaining how it performed relatively well in the centroid distance metric but poorly in the action probe metric.
In Figure 9, the Grad-CAM maps demonstrate that CLIP-Base-32 visualized the simulated and real images completely differently, explaining how it performed poorly in the centroid distance metric despite its low action probe loss. This is a case where the t-SNE plot does not illustrate new information as well about the model’s performance.
@@ V-A Limitations
The relationships between quantitative results and qualitative analysis of the Grad-CAM maps and t-SNE plots were not constant across all observed encoders. The Grad-CAM maps experienced a large amount of variation, even among encoders that performed similarly in both quantitative metrics. Furthermore, several Grad-CAM maps seemed to contradict the encoder’s action probe performance, displaying a significant degree of randomness in the focus of the maps despite performing well in the action probe metric. Additionally, the t-SNE plots rarely exhibited any overlap, even among models that performed well in the centroid distance metric. As such, both qualitative metrics should be further investigated to examine their effectiveness in determining Sim2Real performance.
@@ V-B Future Directions
We propose that the quantitative element of the survey, using centroid distance and linear probing to observe the encoders’ abilities in domain invariance and action inference, could be turned into a database of different encoders and their corresponding performance metrics. Furthermore, this strategy for evaluating encoders could be turned into a score method, accepting different models as inputs and outputting a Sim2Real score calculated through a cost function factoring in centroid distance and action probe performance. We also propose that future research should confirm the findings of this study through real-world experimentation, using encoders to learn transferable visuomotor policies for task completion and testing them in simulated and real environments, comparing performance to confirm their ability to bridge the Sim2Real gap.

### v2

@@ IV-A Analyzing Domain Invariance - Action Inference Plot
We evaluated all 23 encoders with the proposed metrics, Domain Invariance Score (DIS) and Action Score (AS), and present the joint results in Fig. 4. Each point corresponds to a model, with the $x$-axis representing DIS and the $y$-axis representing AS. Encoders closer to the top-left of the plot therefore achieve both strong domain invariance and high action informativeness. Marker color distinguishes CNNs (blue) from transformers (orange), marker shape indicates pre-training type (circle for manipulation-specific data, square for general data), and marker size is proportional to parameter count.
Several clear patterns emerge. First, pre-training data exerts a strong influence: encoders trained on manipulation-relevant datasets consistently attain higher AS values, while encoders trained on generic datasets show wider variance. This trend is especially apparent when comparing ResNet baselines to their R3M counterparts, which are trained with manipulation-focused self-supervision—R3M models consistently achieve higher AS, and in most cases, higher DIS as well. Second, parameter count did not correlate with improved performance: larger encoders did not consistently outperform smaller ones in either metric, suggesting that task relevance of pre-training outweighs raw model capacity. Finally, encoder type plays a role: CNN-based models generally achieve higher DIS than ViTs, implying that convolutional structures preserve greater robustness to the sim–real domain shift.
According to our framework, the highest-performing encoder is MCR (encoder 8) [20], a CNN pre-trained directly on robotic manipulation data, which places near the top-left corner of the plot. This position reflects both strong alignment of sim and real embeddings (high DIS) and effective linear action decoding (high AS), in contrast to encoders such as DinoV2-B that score low on both axes. Together, the joint analysis of DIS and AS offers a quantitative, interpretable benchmark of Sim2Real encoder quality.
IV-B Comparing Quantitative Results To Qualitative Observations
We analyze the framework’s qualitative results using Grad-CAM saliency maps to contextualize the performance of various encoders on the plot. For each encoder, we generated a grid of saliency maps to observe how they relate to the encoder’s quantitative domain invariance and action inference performance. We highlight Fig. 2 and Fig. 3, which contrast the overall worst-performing model (DinoV2-B) and the overall best-performing model (MCR).
Figure 2: Grad-CAM Maps Generated With DinoV2-B
Fig. 2 shows that DinoV2-B (4 in Fig. 4) does not have a clear focus across any of the maps, indicating low action inference performance, which corresponds with its Action Score.
Fig. 3 shows that MCR (8 in Fig. 4) has a sharp focus on the robot end-effectors and the objects in the Grad-CAM maps for both simulated and real images. This finding aligns well with its high performance in both quantitative metrics.
Figure 3: Grad-CAM Maps Generated With MCR
@@ V-A Limitations
While our framework provides a rigorous offline evaluation of encoder transferability, it does not include direct validation on physical robots. As a result, the predictive value of DIS and AS, though theoretically grounded, should ultimately be confirmed in real-world policy deployments. This limitation reflects a broader challenge in Sim2Real research: large-scale robot experiments are costly and difficult to standardize. In addition, although we employed Grad-CAM to provide qualitative interpretability, it is used only as supporting evidence rather than as a primary contribution. Variation in Grad-CAM maps across encoders does not affect the validity of our quantitative metrics, but highlights the inherent noisiness of visual explanation techniques.
@@ V-B Future Directions
Future work can extend this framework in several directions. First, the DIS and AS metrics could form the basis of an open-access benchmark or online database cataloging encoder performance across tasks and domains. Such a resource would allow the robotics community to rapidly compare architectures, training strategies, and pre-training datasets, accelerating progress toward robust visual encoders. Second, these metrics could be incorporated into a reusable software library, enabling practitioners to compute a standardized “Sim2Real score” for any encoder by combining DIS and AS into a single cost function. Finally, integrating offline scores with downstream policy transfer studies on real robots would further validate and calibrate the predictive power of our framework, bridging the gap between mathematical analysis and embodied deployment.